OpenStudy (anonymous):

Find an equation of the tangent line to the graph of F at the given point. MEDAL WILL BE REWARDED! Function: y=x^3+x Point: (-1,-2)

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

did you find the derivative?

OpenStudy (anonymous):

take the derivative \(f'(x)\) then to find the slope at \(-1\) find \(m=f'(1)\) then use the point slope formula

OpenStudy (anonymous):

oops i meant find \(m=f'(-1)\)

OpenStudy (anonymous):

you good from there?

OpenStudy (anonymous):

Im still confused

OpenStudy (anonymous):

\[f(x)=x^3+x\] \[f'(x)=3x^2+1\] \[f'(-1)=3(-1)^2+1=3+1=4\]your slope is \(m=4\)

OpenStudy (anonymous):

the point is \((-1,-2)\) and so using the point slope formula your equation for the line is \[y+2=4(x+1)\] or whatever you get when you solve for \(y\)

OpenStudy (anonymous):

So I just plug in the point values into the equation?

OpenStudy (anonymous):

not sure what you mean exactly \(f'(x)\) is the formula for the slope the actual slope is a number, when you get when you compute \(f'(-1)\) in this example

OpenStudy (anonymous):

Oh, okay, now I see! Thank you so much!!!!

OpenStudy (anonymous):

slope is \(f'(-1)=4\)and point is \((-1,-2)\) and point - slope formula gives equation of the line yw

OpenStudy (anonymous):

Oh okay I completely understand this now

OpenStudy (anonymous):

Thank you for making it easier